OpenStudy (anonymous):

I GIVE MEDALS!!!!!!! Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. A.) y − 6 = −3x − 3 B.) y − 6 = −3(x + 1) C.) y = −3x + 3 D.) 3x + y = 3

OpenStudy (anonymous):

point slope is y-y1=m(x-x1) y1 is the y coordinate x1 is the x coordinate m is the slope y-6=-3(x-(-1)) or y-6=-3(x+6)

OpenStudy (anonymous):

so C?

OpenStudy (anonymous):

no b it says to give the point-slope not the slope-intercept

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

if it said to give slope-intercept then it would be c

OpenStudy (anonymous):

i get it now...thanx soo much

OpenStudy (anonymous):

yep